# Multi-scale Restoration of Missing Data in Optical Time-series Images with Masked Spatial-Temporal Attention Network

### v1

@@ IV-B1 Simulated ETM+ SLC-off Restoration Experiment
For the land satellite ETM+ SLC-Off problem, we conducted simulated experiments with synthesized missing data in research areas S1 and S2, as illustrated in Fig. 7-8. In these figures, (a) and (b) represent the original image and the simulated SLC-off image, respectively, while (c)-(f) show the simulated recovery results using four methods—LLHM, WLR, STS-CNN, and the proposed MS2TAN. To provide a clearer view of the recovery results, we present magnified details next to each reconstruction. As shown in Fig. 7-8 (c)-(e), all compared methods generate discontinuous fine features to some extent. This is because the temporal auxiliary images cannot fully cover the missing regions, thus requiring LLHM and WLR algorithms to rely on LPRM to fill the remaining gaps, resulting in highly blurred noise bands in the reconstructed results. Although the STS-CNN model based on an end-to-end strategy can repair these gaps, it still cannot accurately restore the original features. In contrast, the proposed MS2TAN, utilizing a Masked Spatial-Temporal Attention mechanism and integrating a “Pixel-Structure-Perception” Multi-Objective Optimization approach, achieves better restoration of texture and structural details, as shown in Fig. 7-8 (f). Moreover, this method outperforms state-of-the-art methods in the quantitative evaluations presented in Table II, validating its strong capability in spatio-temporal integrated data utilization.
Figure 9: Experimental results of Simulated Thick Cloud Removal based on Research Area S3. (a) Ground truth (October 17, 2003). (b) Simulated Thick Cloud. (c) LLHM. (d) WLR. (e) STS-CNN. (f) Ours.
Figure 10: Experimental results of Simulated Thick Cloud Removal based on Research Area S4. (a) Ground truth (January 22, 2014). (b) Simulated Thick Cloud. (c) LLHM. (d) WLR. (e) STS-CNN. (f) Ours.
TABLE II: Quantitative Evaluation Result of the Simulated SLC-off Restoration in Research Area S1 and S2
  Dataset Area S1 Area S2
Metric mPSNR($\uparrow$) mSSIM($\uparrow$) mPSNR($\uparrow$) mSSIM($\uparrow$)
  LLHM [21] 32.424 0.8947 32.908 0.9049
WLR [24] 34.374 0.9179 34.498 0.9214
STS-CNN [37] 37.118 0.9446 37.761 0.9479
$\textbf{MS}^{2}\textbf{TAN}$ (Ours) 38.255 0.9548 38.963 0.9555
IV-B2 Simulated Thick Cloud Removal Experiment
Similar to the simulated experiments of ETM+ SLC-off, we simulated thick cloud removal tasks in research areas S3 and S4, and reconstructed the data using the same four methods, resulting in the outcomes shown in Fig. 9-10. Among them, LLHM exhibits severe texture discontinuities, while WLR produces numerous noise points. This is because these simple models struggle to handle highly complex nonlinear relationships between different times. In comparison, the STS-CNN model can more accurately reconstruct the texture and structural features of missing parts, but exhibits noticeable spectral transition phenomena at the edges of missing value regions. On the other hand, the proposed MS2TAN method introduces missing value hint tensors into the learning process, enhancing the model’s capability to handle irregularly distributed missing values. Ultimately, it performs well in maintaining consistency in texture, structure, and spectral properties, achieving the best metrics in the quantitative evaluation presented in Table III.
TABLE III: Quantitative Evaluation Result of the Simulated Thick Cloud Removal in Research Area S3 and S4
  Dataset Area S3 Area S4
Metric mPSNR($\uparrow$) mSSIM($\uparrow$) mPSNR($\uparrow$) mSSIM($\uparrow$)
  LLHM [21] 34.775 0.9439 34.714 0.9370
WLR [24] 39.003 0.9639 38.261 0.9614
STS-CNN [37] 39.515 0.9711 38.753 0.9662
$\textbf{MS}^{2}\textbf{TAN}$ (Ours) 39.802 0.9753 39.272 0.9693
Figure 11: Experimental results of Real ETM+ SLC-off Restoration based on Research Area S5. (a) SLC-off Image 1 (October 23, 2011). (b) SLC-off Image 2 (November 8, 2011). (c) LLHM+LPRM. (d) WLR+LPRM. (e) STS-CNN. (f) Ours.
Figure 12: Experimental results of Real Thick Cloud Removal based on Research Area S6. (a) Image with clouds (August 5, 2007). (b) Temporal image without clouds (July 20, 2007). (c) LLHM. (d) WLR. (e) STS-CNN. (f) Ours.
IV-C Real Data Experiment
IV-C1 SLC-off Restoration Experiment
The experimental results of real SLC-off images in research area S5 are shown in Fig. 11. In this figure, (a) and (b) depict two ETM+ SLC-off images observed on October 23, 2011, and November 8, 2011, respectively, while (c)-(h) display the outputs of four algorithms. Similar to the simulated SLC-off reconstruction experiment, single temporal auxiliary images cannot fully cover the gaps. Thus, after LLHM and WLR processing, the remaining gaps are filled using the LPRM algorithm. STS-CNN can directly repair the entire image but exhibits discontinuous texture features similar to the simulated experiment. Thanks to the “Pixel-Structure-Perception” Multi-Objective Joint Optimization method, the proposed MS2TAN method preserves more structural details and produces more consistent reconstruction results.
IV-C2 Thick Cloud Removal Experiment
The experimental results of thick cloud coverage reconstruction in region S6 are shown in Fig. 12. In Fig. 12 (a), an image containing thick clouds is presented, while (b) depicts the temporal auxiliary image. Four algorithms identical to those used in the SLC-off reconstruction experiment were employed to reconstruct the images, yielding results displayed in (c)-(f). It can be observed in Fig. 12 (a)-(b) that for extensive cloud coverage and cloud shadows, the results of LLHM and WLR models suffer from significant loss of texture and structural features, presenting as blurry patches. This is attributed to the inconsistency between the reconstructed texture details in the cloud region and the surrounding non-cloud areas, which the aforementioned models fail to fit due to their incapability to capture the nonlinear relationship between different temporal data. In comparison, the STS-CNN method exhibits better performance in texture restoration, yet still shows noticeable spectral discontinuities at the boundaries of missing regions. On the other hand, the proposed MS2TAN method fully preserves texture and structural details with less spectral distortion compared to STS-CNN, resulting in high-quality thick cloud removal images.
IV-D Validation Study
IV-D1 Masked Spatial-Temporal Attention
To validate the effectiveness of the proposed Masked Spatial-Temporal Attention (MSTA) mechanism, we compared models with and without the adoption of this mechanism, and plotted their mSSIM iteration curves on the validation set, as shown in Fig. 13. It can be observed that with the assistance of the attention mask mechanism, the model exhibits faster convergence speed and higher fitting accuracy. This confirms that the attention mask mechanism effectively enhances the model’s feature extraction capability and training speed.
Figure 13: Ablation experiments for the proposed MSTA mechanism. The blue and red curves respectively depict the variations of mSSIM with training rounds for models with and without the adoption of the attention mask mechanism in Area S1.
IV-D2 Multi-scale Restoration
The MS2TAN employs a multi-scale restoration strategy to process input images. To validate its effectiveness, we take the SLC-off reconstruction task as an example and compare the intermediate outputs of each layer. The results are shown in Fig. 14. In the figure, the missing regions of the input image are filled with zeros. After Coarse Embedding, MFE, and Unembedding, a coarse repair result is obtained. This result reconstructs most of the texture and structural features and balances the significant distribution differences between the missing regions and other areas in the input. However, there are still spectral discontinuities at the edges of the missing regions. After several layers of restoration, the final refined repair result is obtained. This result refines the previous intermediate results, with clearer textures, consistent spectra, and smoother integration of the repair area boundary with other areas.
Figure 14: Ablation experiments for multi-scale restoration. (a)-(c) represent observations from three different times, processed in parallel by MS2TAN after composing the time-series remote sensing images. From top to bottom, they represent: input with missing values, coarse result, and fine result.
IV-D3 Gains from long time-series image inputs
A highlight of the proposed model is its capability to utilize longer time-series images for assistance in restoration, leading to more accurate reconstruction results. We designed experiments to compare different lengths of time-series inputs to quantitatively evaluate the accuracy gains from longer time-series inputs. The experimental results are shown in Fig. 15. From Fig. 15, it can be observed that as the length of the input time series increases from 2, there is a significant improvement in mPSNR and mSSIM metrics. However, when the time series length reaches 5 or longer, the change in accuracy becomes less significant. Therefore, longer time-series inputs can significantly improve the accuracy of missing value reconstruction, but there is a certain limit. Hence, we can find a balance point between reconstruction accuracy and algorithm flexibility.
Figure 15: Analysis of the input time-series length. The left and right tables record the variations of mPSNR and mSSIM with the length of the time-series, respectively, and distinguish the results of four research areas with different colors.
@@ V Conclusion
In this paper, we proposed a novel method for time-series remote sensing image reconstruction based on a Multi-scale Masked Spatial-Temporal Attention Network. The proposed method reconstructs various missing value scenarios, including the Landsat ETM+ SLC-off problem and removal of thick cloud coverage in remote sensing images. Compared to existing methods, the proposed method improves the efficiency of spatio-temporal information utilization, enabling the mutual assistance of multiple degraded time-series images for reconstruction and achieving higher restoration accuracy through more auxiliary information. Additionally, the application of the MSTA mechanism and Multi-Objective Joint Optimization further enhances the texture and structural consistency of the reconstruction results. Experimental results demonstrate significant improvements in visual effects and restoration accuracy compared to mainstream methods.
Although the proposed method performs well in reconstructing the ETM+ SLC-off problem and thick cloud removal, there are still some inevitable limitations. For example, the model needs to simulate various shapes of cloud-like missing regions in the dataset and learn the characteristics of missing value distribution extraction sufficiently to achieve better cloud removal effects, which requires more training time for the model. In our future research, we can incorporate network components pre-trained on large-scale datasets into the model to improve its convergence speed.
@@ Acknowledgments
This work was supported in part by the National Key R&D Program of China (Project No. 2022YFC3800700); in part by the National Natural Science Foundation of China (No. U21A2013).

### v2

@@ IV-C1 Comparison under Dual-Temporal Inputs
We conducted comparison experiments with the DISO algorithms using synthetic data gaps in the Landsat and EarthNet2021 datasets. As illustrated in Fig. 5, (a) and (b) represent the two images with data gaps used as input, while (c)-(g) sequentially display the recovery results obtained from five methods, with (h) showing the ground truth of the target image. As depicted in (c)-(f), all compared methods to some extent produced discontinuous fine features. This is because the time-assisted images were unable to completely cover the missing regions, leading the LLHM, and WLR to rely on LRPM (Laplacian Prior Regularization Method [24]) to fill in the remaining gaps, resulting in very blurry noise bands in the reconstruction results. Although the end-to-end strategy-based STS-CNN model can repair these blanks, it fails to accurately restore the original features. In contrast, the proposed MS2TAN, integrating a “Pixel-Structure-Perception” Multi-Objective Joint Optimization approach, achieves superior texture and structural detail recovery, as shown in (g). In the quantitative evaluation shown in Table I, MS2TAN outperforms other comparative methods across all evaluation metrics. Compared to the previous SOTA method STS-CNN, MS2TAN exhibits a 10.99%/6.86% (in two datasets respectively) reduction in MAE and a 0.75 dB/0.53 dB increase in PSNR, showcasing its robust capability in utilizing spatiotemporal data.
IV-C2 Comparison under Multi-Temporal Inputs
Figure 7: Visual evaluation of restoration results for real ETM+ SLC-off images from the Landsat dataset, with dead pixels shown in black. The first and second rows show the observed sequence and the restoration results of the MS2TAN, respectively.
Figure 8: Visual evaluation of restoration results for real Thick Cloud Cover images in the EarthNet2021 dataset, with clouds shown in white. The first and second rows show the observed sequence and the restoration results of the MS2TAN, respectively.
Similar to the dual-temporal input experiment, we conducted comparative experiments with the MIMO algorithms under multi-temporal input conditions. Visual comparative results are illustrated in Fig. 6, displaying, from top to bottom: the input image sequence, the restoration outcomes of five methods, and the ground truth. Since traditional linear methods rely on time-series information, they cannot restore regions in the input sequence where all time steps are missing. In contrast, U-TILISE can recover data gaps of any shape, but still exhibits significant color shifts in irregularly shaped gaps, reducing the quality of the results, as shown in the first three columns of Fig. 6. Benefiting from the Masked Spatial-Temporal Attention mechanism, MS2TAN can effectively handle the edges of missing regions, achieving more harmonious seams. Notably, the Multi-Objective Optimization makes MS2TAN perform better in restoring structural and textural details in the missing regions. In some extreme cases (e.g., the upper-left region of the first 8 columns), due to a lack of sufficient information, MS2TAN may not accurately reproduce the actual changes but can still output coherent images with natural color transitions. Compared to the U-TILISE, MS2TAN exhibits a 9.76%/9.30% reduction in MAE, a 0.56 dB/0.62 dB increase in PSNR, and achieves higher visual consistency.
IV-C3 Restoration of the Real Data Gaps
To evaluate the generalization performance of the proposed MS2TAN, we applied pre-trained model weights to real ETM+ SLC-off sequences and thick cloud-covered sequences. Due to the lack of ground truth reflectance data, the assessment was conducted based solely on visual inspection. Visualization of the restoration results is provided in Figs. 7 and 8. In the restored ETM+ SLC-off images, the roads and rivers, which are divided by striped missing areas, are rationally connected in the restoration results with the original structure maintained. Obvious ground surface changes brought by long time intervals are also taken into account. MS2TAN utilizes Masked Spatial-Temporal Attention to learn the distribution of missing data, thereby capturing temporal evolution patterns from neighboring frames and fitting spatial textures, achieving high-precision data restoration. For the thick cloud-covered images, the method accurately restored all missing data, including both large-scale and fragmented gaps. Based on the assistance of long time-series, continuous large-area data gaps can also be repaired reasonably well and keep the original evolutionary trend. Irregularly shaped data gaps (e.g., the images in the sixth column in Fig. 8) are repaired to blend in with the spatial domain, preserving high textural and structural coherence. By applying Masked Spatial-Temporal Attention and Multi-Objective Joint Optimization, MS2TAN exhibits excellent performance in handling complex and large-scale data gaps, which often present challenges in traditional data recovery methods, as evident in the detailed visual evaluation.
IV-D Validation Studies
To validate the effectiveness of the current architecture, we experimented with some variants of MS2TAN and obtained quantitative results as shown in Table III. We discuss each of the mechanisms used below.
TABLE III: Quantitative Evaluation of Different MS2TAN Variants that derived from whether Separated Spatial-Temporal Attention (Sep. Attn), Attention Mask Mechanism (Attn Mask), and Multi-scale Restoration Strategy (Multi-Scale) are enabled. Statistics information and evaluation metrics are provided on the right. All experiments are performed under Multi-Temporal Inputs.
       Features Statistics Landsat EarthNet2021
Sep. Attn Attn Mask Multi-Scale Params (M) FLOPs (G) $\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM
  4.33 8.756 0.0074 0.96 39.01 0.9552 0.0078 1.71 38.91 0.9728
  4.33 14.626 0.0073 0.95 39.08 0.9560 0.0083 1.81 38.53 0.9714
4.33 8.756 0.0096 1.18 37.47 0.9384 0.0112 2.38 36.37 0.9609
6.61 10.130 0.0086 1.07 38.09 0.9459 0.0095 2.05 37.60 0.9672
6.61 16.473 0.0101 1.25 37.24 0.9356 0.0127 2.73 35.19 0.9538
IV-D1 Separated Spatial-Temporal Attention
The original ViT, after embedding images into token sequences, feeds the entire sequence into a multi-head self-attention module. However, the computational complexity of self-attention is proportional to the square of the sequence length. For a sequence of remote sensing images ($t$ images, $n$ patches per image), the computational complexity is as high as $O\left(t^{2}\times n^{2}\right)$. Whereas, separating temporal attention from spatial attention and computing them sequentially reduces the complexity to $O\left(t^{2}+n^{2}\right)$. Although the perceptual field of the separated spatial-temporal attention is reduced in a single operation, we alternate multiple temporal and spatial attention modules to spread the perceptual field to the whole spatiotemporal range. Meanwhile, the extension of the perceptual field along the temporal axis and the spatial plane matches the characteristics of the time-series remote sensing images, which have small displacement in time and large spatial correlation, and can fully integrate the spatial and temporal information. As shown in the first two rows of Table III, the use of separated spatial-temporal attention reduces the computational effort by 40.1% with almost the same accuracy in the Landsat dataset, while the accuracy in EarthNet2021 is slightly improved.
Figure 9: Visualization of Temporal Attention Distribution in the first MSTA unit. We present the attention scores for repairing the 3rd frame (highlighted with a red box in row 1) of the sequence. It is observed that, after incorporating the mask, the temporal attention is no longer restricted to the vicinity of the target time but instead focuses on a broader range of valid pixels. From top to bottom: Input (row 1), w/o Mask (rows 2-5), and w/ Mask (rows 6-9).
Figure 10: Visualisation of Spatial Attention Distribution in the first MSTA unit. We show the average attention scores for imputing each patch. With the mask, spatial attention focuses on the edges of the valid pixels, extracting textural and structural features, rather than spreading around as it would be without the mask. From top to bottom: Input (row 1), w/o Mask (rows 2-3), and w/ Mask (rows 4-5).
Figure 11: Ablation Experiment Results for Optimal $C_{\text{max}}$ Value in MSTA. The horizontal axis represents different $C_{\text{max}}$ values, and the vertical axis represents the SSIM in the EarthNet2021 dataset.
IV-D2 Attention Mask in MSTA
To verify the effectiveness of the attention mask in MSTA, we visualized the distribution of attention. Fig. 9 shows the distribution of temporal attention with and without the attention mask, while Fig. 10 illustrates the spatial attention distribution. Without the mask, certain attention heads focus excessively on patches with missing data, while others focus excessively on their own patch, providing no additional information to impute missing values and leading to unnecessary computational costs. With the mask mechanism in place, the attention heads no longer focus on missing-data patches or themselves, but instead extract information from spatiotemporal domains, obtaining a larger effective receptive field to achieve higher Imputation accuracy. Within the MSTA framework, the maximum allowable missing rate, $C_{\text{max}}$, is a crucial hyperparameter. Setting it too high can cause patches with missing data to distort token distribution, while a too-low value may result in information loss. Consequently, we conducted ablation experiments on different values of $C_{\text{max}}$, with the results shown in Fig. 11. The findings indicate that an optimal $C_{\text{max}}$ yields the best reconstruction accuracy, and thus $C_{\text{max}}$ is set to 0.5 for subsequent experiments in this paper.
IV-D3 Multi-scale Restoration
Figure 12: Visualisation of intermediate results for multi-scale restoration. The values above indicate the MAE compared to the ground truth. (a) Input. (b)-(d) Intermediate outputs at each scale. (e) Ground truth.
MS2TAN employs a multi-scale restoration strategy to refine the restoration results layer by layer. To verify its effectiveness, we take the SLC-off reconstruction task as an example and compare the intermediate outputs of each layer in Fig. 12. With the gradual deepening of the restoration stages, the restoration results are gradually refined to better capture the details and structural features in the image. Specifically, the coarse repair stage addresses global textures and structural issues, while the fine repair stage further optimizes local details and spectral consistency. However, despite the effectiveness of the multi-scale approach, introducing too many scale levels leads to diminishing returns, with further improvements becoming negligible or even introducing overfitting and increased computational costs. Therefore, selecting an appropriate number of scale levels is crucial.
IV-D4 Multi-Objective Joint Optimization
TABLE IV: Ablation Analysis for Multi-Objective Joint Optimization, where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ denote the weights of the pixel-wise loss, structural loss and perceptual loss respectively.
         Loss weights       EarthNet2021
  $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$downarrow MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM
  1 - - 0.0120 2.55 35.76 0.9572
0.9 0.1 - 0.0091 1.97 37.94 0.9686
0.9 - 0.1 0.0102 2.20 37.09 0.9645
- 0.5 0.5 0.0210 4.43 31.72 0.9241
  0.9 0.05 0.05 0.0078 1.71 38.91 0.9728
In Multi-Objective Joint Optimization, the weights assigned to each loss function are critical. To investigate this, we conducted an ablation study on the allocation of loss weights. Table IV presents different weight configurations for pixel-wise loss ($\lambda_{1}$), structural loss ($\lambda_{2}$), and perceptual loss ($\lambda_{3}$), along with their corresponding results on the EarthNet2021 dataset. The findings show that combining pixel-wise loss with either structural or perceptual loss leads to significant performance improvements (cf. rows 2-3), with the best results achieved when both are combined. However, when only structural and perceptual losses are applied, the model struggles to converge, resulting in poor performance (cf. row 4). Therefore, pixel-wise loss plays a dominant role in multi-objective optimization, and its combination with structural and perceptual losses is essential for achieving high restoration accuracy and visual consistency.
TABLE V: Hyperparameter settings and Evaluation of MS2TAN with different sizes. The hyperparameters include the restoration scales number ($S$), patch size ($P$), embedding dimensions ($d_{\text{emb}}$), attention heads number ($h$), dimensions of vectors Q, K and V in self-attention ($d_{\text{qkv}}$), and the MSTA units number ($L$), where bracketed lists indicate the hyperparameters for each scale.
  Model Hyperparameters Statistics EarthNet2021
$S$ $P$ $d_{\text{emb}}$ $h$ $d_{\text{qkv}}$ $L$ Params (M) FLOPs (G) $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM
  MS2TAN 3 (12, 10, 8) (256, 192, 128) (8, 6, 4) (32, 32, 32) (2, 2, 2) 4.33 8.756 0.0078 1.71 38.91 0.9728
  MS2TAN-L 3 (12, 10, 8) (384, 256, 192) (8, 6, 4) (48, 48, 48) (4, 4, 4) 17.04 34.363 0.0076 1.67 39.03 0.9733
MS2TAN-S 2 (12, 10) (192, 128) (8, 6) (24, 24) (2, 2) 2.07 3.567 0.0089 1.92 38.07 0.9692
IV-D5 Balance between Effectiveness and Efficiency
Scalability is a key strength of ViT-based models, and MS2TAN is no exception. To strike a balance between effectiveness and efficiency, we adjusted several critical hyperparameters to obtain two MS2TAN variants and evaluated them on EarthNet2021 dataset (see Table V). The experimental results reveal that when the model is too small (MS2TAN-S), performance significantly degrades; conversely, when the model is too large (MS2TAN-L), there is only a slight performance gain, while the computational cost increases drastically, making the approach less efficient. Overall, the current configuration of parameters and computational load represents an optimal “sweet spot” where good performance is maintained without excessive computational overhead.
@@ V Conclusion
This paper presents a novel approach to reconstruct time-series remote sensing images using the Multi-Scale Masked Spatial-Temporal Attention Network (MS2TAN). MS2TAN takes an image sequence with arbitrary data gaps as input and generates a clear and complete time-series. Compared with the existing methods, the proposed Masked Spatial-Temporal Attention (MSTA) mechanism improves the efficiency of spatiotemporal information utilization, reduces the high computational cost of the original ViT, and obtains higher restoration accuracy. In addition, the application of the Multi-scale Restoration strategy and Multi-Objective Joint Optimization further enhances the texture and structural consistency of the reconstructed images. MS2TAN achieves higher imputation accuracy than the mainstream methods in quantitative experiments and exhibits better visual effects in eliminating real data gaps. Ablation experiments also confirm the contributions of primary innovations to the results.
Despite the effectiveness of the proposed method in addressing ETM+ SLC-off and thick cloud cover, there are some unavoidable limitations. For instance, the model needs to simulate various shapes of data gaps in the dataset to adequately learn to extract the features of the missing value distribution in order to achieve better restoration results, which leads to the model needing more time for training. Therefore, future research could incorporate network components pre-trained on large-scale datasets to improve model convergence speed and generalization performance.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (No. 42471505); in part by the National Key R&D Program of China (Project No. 2022YFC3800700).
Table VI details the statistics of the dataset we used, including key information such as data source, resolution, and number of sequences.
TABLE VI: Details of the Landsat and EarthNet2021 Datasets.
  Landsat       EarthNet2021
        Data source Landsat-5/7/8       Sentinel-2
Restoration task SLC-off & Thick Cloud Thick Cloud
Spatial scope China & USA Europe
Temporal scope 2001 - 2011 2016 - 2020
        Spectral bands 6 bands       4 bands
Image size 120 $\times$ 120 128 $\times$ 128
Spatial Resolution 30-m 20-m
Temporal length 10 30
Temporal interval average 30 days constant 5 days
        Number of samples 11913       28123
– Train 7625 18955
– Valid 1905 4949
– Test 2383 4219